OpenStudy (anonymous):

Find the equilibrium concentrations of A, B, and C for a=1, b=1, and c=2. Assume that the initial concentrations of A and B are each 1.0 M and that no product is present at the beginning of the reaction. aA(g)+bB(g)⇌cC(g), Kc = 4.0

OpenStudy (alphadxg):

A B C Initial Concentration 1.0M 1.0M 0.0M Change in Concentration -x -x +2x Equilibrium Concentration 1-x 1-x 2x 4.0 = C^2/(A)(B) C=2x, A=B=1-x 4.0=[(2x)^2]/[(1-x)^2] 4(1-x)^2=4x^2 4-4x^2=4x^2 8x^2=4 x^2=4/8=1/2 x=0.707 Equilibrium Concentrations A: 1-x=1-0.707=0.293 M B: 1-x=0.293 M C: x=1.414 M